OpenStudy (anonymous):

Prove that the function f:[1,∞)→[0,∞) defined by f(x) = sqrt(x-1) is continuous at x=10. @ganeshie8 @myininaya @UnkleRhaukus @RadEn @robtobey @chmvijay @Luigi0210 @eliassaab @genius12 @Vincent-Lyon.Fr @Preetha @Kainui @ehuman @wolfe8 @lucaz @INeedHelpPlease? @khadeeja @thadyoung @A_clan @tester97 @Andras @Confusionist @linh412986 @BlackLabel @leozap1 @JMark @Microrobot @zacharyf @Rubio101 @NaomiBell1997 @MeganEdward @theballer225 @link,zoro @Goldenmoon @CrayolaCrayon_

OpenStudy (anonymous):

Wow @doggy can I have your account or something your account is a WIN.

OpenStudy (anonymous):

f(x) is continuous at x = 10 if the limit of the function at x = 10 equals f(10). We show this as follows: \[\bf \large \lim_{x \rightarrow 10} \sqrt{x-1}=3=f(10)\]Hence the function is continuous at x = 10.